SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (2nd Jan, 2007)

Attendees: Debbie Bard, Al Eisner, Selina Li, Su Dong.

Operations

  • Cosmics with B field will be taken this weekend - we should probably take some good data first with the boards in IR2 (BLT and GLT) and then swap in the spares to check they're alright.
  • New proms in the BLT and GLT - need to update web pages with what's written on the stickers so there's no confusion.
  • Still problems running L1TNtuple on the cosmics run 69139.
  • Still need to check with IFR people if timing jumper was switched to LST.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.